# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

### v1

@@ 1. $\bf(1)$}
$\rho_{\infty}(\xi,\tilde{\xi}):=\sup_{t\in[0,T]}\left\{e^{-\lambda t}|\xi_{t}-\tilde{\xi}_{t}|\right\},\quad\xi,\tilde{\xi}\in\mathcal{D}([0,T];\mathbb{R}^{d}),\quad\lambda>0.$
@@
The following lemma focuses on the completeness of $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, which is essentially due to Shao et al. [40, Lemma 2.3 ]. We include a complete proof for readers’ convenience.
Lemma 2.1.
The metric space $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ is complete for every $p\geq 1$.
Proof.
Let $(X^{(n)})_{n\geq 1}$ be a Cauchy sequence in $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, i.e.,
$\bm{\rho}_{p,\infty}(X^{(n)},X^{(m)})^{p}\to 0\quad\text{as }n,m\to\infty.$
Take a subsequence $(X^{(n_{k})})$ such that
$\bm{\rho}_{p,\infty}(X^{(n_{k})},X^{(n_{k+1})})<\frac{1}{2^{k}},\quad k\geq 1,$
which implies
$\mathbb{P}\left(\rho_{\infty}(X^{(n_{k})},X^{(n_{k+1})})\geq\frac{1}{k^{2}}\right)\leq\frac{k^{2p}}{2^{kp}},\quad k\geq 1,$
Since $\sum_{k\geq 1}\frac{k^{2p}}{2^{kp}}<\infty$, the Borel-Cantelli lemma yields
$\mathbb{P}\left(\rho_{\infty}(X^{(n_{k})},X^{(n_{k+1})})\geq\frac{1}{k^{2}},\text{ i.o.}\right)=0.$
Thus, there exists a measurable set $\Xi$ with $\mathbb{P}(\Xi)=1$ such that for every $\omega\in\Xi$,
$\sum_{k=1}^{\infty}\rho_{\infty}(X^{(n_{k})}(\omega),X^{(n_{k+1})}(\omega))<\infty.$
By the completeness of $\mathcal{D}([0,T];\mathbb{R}^{d})$ with respect to $\rho_{\infty}$, for every $\omega\in\Xi$, there exists $\tilde{X}(\omega)\in\mathcal{D}([0,T];\mathbb{R}^{d})$ such that
$\lim_{k\to\infty}\rho_{\infty}\bigl{(}X^{(n_{k})}(\omega),\,\tilde{X}(\omega)\bigr{)}=0,\quad\omega\in\Xi.$
We now aim to prove that $\tilde{X}\in\mathscr{Y}_{T}(p)$ and that
$\bm{\rho}_{p,\infty}\big{(}X^{(n_{k})},\tilde{X}\big{)}\xrightarrow{\;k\to\infty\;}0.$
We first show that the Cauchy sequence $(X^{(n)})_{n\geq 1}$ is uniformly $p$-integrable. Since $(X^{(n)})$ is Cauchy in $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, there exists some fixed index $n_{0}$ and a constant $C<\infty$ such that
$\sup_{n\geq 1}\bm{\rho}_{p,\infty}\big{(}X^{(n)},X^{(n_{0})}\big{)}^{p}=\sup_{n\geq 1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{0})}_{t}|^{p}\Big{]}\leq C.$
Using the elementary inequality $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$, we obtain for every $n\geq 1$
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}|^{p}\Big{]}$ $\leq 2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{0})}_{t}|^{p}\Big{]}+2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{0})}_{t}|^{p}\Big{]}$
$\leq 2^{p-1}C+2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{0})}_{t}|^{p}\Big{]}=:C_{1}<\infty.$
Hence $\sup_{n\geq 1}\mathbb{E}\big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}|^{p}\big{]}\leq C_{1}<\infty$. Now consider the limit process $\tilde{X}$. By Fatou’s lemma and the Cauchy property of $X^{(n_{k})}\to\tilde{X}$,
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|\tilde{X}_{t}|^{p}\Big{]}\leq\liminf_{k\to\infty}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{k})}_{t}|^{p}\Big{]}\leq C_{1}<\infty.$
[…]
$\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]\leq e^{\lambda T}\,\mathbb{E}\Big{[}\sup_{s\in[0,T]}e^{-\lambda s}|\tilde{X}_{s}|^{p}\Big{]}.$
Since the right-hand side has already been shown to be finite (bounded by $C_{1}$), we conclude
$\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]\leq e^{\lambda T}C_{1}<\infty.$
Thus we have $\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]<\infty$.
Note that for every $\omega\in\Xi$, the path $\tilde{X}(\omega)$ belongs to $\mathcal{D}([0,T];\mathbb{R}^{d})$. Thus, it remains only to verify that $\tilde{X}_{t}$ is $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$. Let $\tilde{\Gamma}$ denote the law of $\tilde{X}$ as a random element of $\mathcal{D}([0,T];\mathbb{R}^{d})$. We introduce the set
$\mathcal{F}_{\tilde{\Gamma}}:=\{\,t\in[0,T]\mid\tilde{\Gamma}(J_{t})=0\,\},$
where
$J_{t}:=\{\,X\in\mathcal{D}([0,T];\mathbb{R}^{d}):X_{t}\neq X_{t-}\,\}$
[…]
$\lim_{k\to\infty}|X_{t}^{(n_{k})}-\tilde{X}_{t}|=0,\quad\mathbb{P}\text{-a.e.},\forall t\in\mathcal{F}_{\tilde{\Gamma}}$.
Since $X_{t}^{(n_{k})}$ is $\mathcal{F}_{t}$-adapted, $\tilde{X}_{t}$ is $\mathcal{F}_{t}$-adapted for $t\in\mathcal{F}_{\tilde{\Gamma}}$. Due to right‐continuity of the paths, for any $t\in(0,T)$, we choose a sequence $t_{n}\downarrow t$ such that $\tilde{X}_{t_{n}}\;\xrightarrow{\;n\to\infty\;}\;\tilde{X}_{t}\quad\mathbb{P}\text{-a.s.}.$ Since $\mathcal{F}_{t+}=\mathcal{F}_{t}$, it follows that $\tilde{X}_{t}\;\in\;\mathcal{F}_{t}.$ Hence, $\tilde{X}\in\mathscr{Y}_{T}(p)$. By Fatou’s lemma and the Cauchy property,
[…]
$=$ $\lim_{n\to\infty}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-\lim_{l\to\infty}X^{(n_{l})}_{t}|^{p}\bigg{]}^{1/p}$
$\leq$ $\lim_{n\to\infty}\mathop{\underline{\lim}}\limits_{l\to\infty}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{l})}_{t}|^{p}\bigg{]}^{1/p}=0.$
Thus, $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ is complete. ∎
@@ 1. $\bf(2)$}
almost surely for any $t\in[0,T],$ with $Y_{0}^{\mu}=X_{0}$. According to [37, Theorem 2.1], the classical SDE (2.1) admits a unique solution with $\sup_{t\in[0,T]}\mathbb{E}\bigl{[}\lvert Y_{t}^{\mu}\rvert^{2}\bigr{]}<\infty.$ Moreover, we have $(Y_{t}^{\mu})_{t\in[0,T]}\in\mathscr{Y}_{T}(2)$.
@@ 1. $\bf(3)$}
We show that $\Phi$ is a strict contraction in $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$. For another $\tilde{\eta}\in\mathscr{Y}_{T}(2)$, let $(Y_{t}^{\nu})$ solve SDE (2.1) by replacing $\mu_{t}=\mathcal{L}(\eta_{t}|\mathcal{F}_{t}^{0})$ there with $\nu_{t}=\mathcal{L}(\tilde{\eta}_{t}|\mathcal{F}_{t}^{0})$ for $t\in[0,T]$. Applying Itô’s formula to $|Y_{t}^{\mu}-Y_{t}^{\nu}|^{2}$, taking the supremum over $t\in[0,T]$, taking expectation and using the Assumption (A2), we get
@@
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big{(}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\delta_{s}(z)|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big{)}\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
@@
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big{(}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\delta_{s}(z)|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big{)}\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
Applying the Burkholder-Davis-Gundy (BDG) inequality, Jensen’s inequality and using the Assumption (A7) to each stochastic integral, we obtain
$\mathbb{E}\left[\sup_{0\leq r\leq t}\left|\int_{0}^{r}\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}\rangle\right|\right]$
@@
$\quad\leq K_{1}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
where $K_{1}:=K_{1}(C_{1},\tilde{L}_{1},\tilde{L}_{2})$ and similarly for the $W^{0}$ term and the jump term with $K_{2},K_{3}$ respectively. Therefore,

### v2

@@ 1. $\bf(1)$}
$\rho_{\infty}(\xi,\tilde{\xi}):=\sup_{t\in[0,T]}\left\{|\xi_{t}-\tilde{\xi}_{t}|\right\},\quad\xi,\tilde{\xi}\in\mathcal{D}([0,T];\mathbb{R}^{d}).$
@@
The completeness of $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ follows essentially from Lemma 2.3 in Shao et al. [31].
@@ 1. $\bf(2)$}
almost surely for any $t\in[0,T],$ with $Y_{0}^{\mu}=X_{0}$. According to [28, Theorem 2.1], the classical SDE (2.1) admits a unique solution with $\sup_{t\in[0,T]}\mathbb{E}\bigl[\lvert Y_{t}^{\mu}\rvert^{2}\bigr]<\infty.$ Moreover, we have $(Y_{t}^{\mu})_{t\in[0,T]}\in\mathscr{Y}_{T}(2)$.
@@ 1. $\bf(3)$}
We show that $\Phi$ is a strict contraction in $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$. For another $\tilde{\eta}\in\mathscr{Y}_{T}(2)$, let $(Y_{t}^{\nu})_{t\in[0,T]}$ solve SDE (2.1) by replacing $\mu_{t}=\mathcal{L}(\eta_{t}|\mathcal{F}_{t}^{0})$ there with $\nu_{t}=\mathcal{L}(\tilde{\eta}_{t}|\mathcal{F}_{t}^{0})$ for $t\in[0,T]$. Applying Itô’s formula to $|Y_{t}^{\mu}-Y_{t}^{\nu}|^{2}$, taking the supremum over $0,t]\subset[0,T]$ and expectation, and using the Assumption (A2), we get
@@
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
@@
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right].$
Moreover, due to
$\mathbb{E}[|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}]\leq\mathbb{E}\Big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big],\qquad\mathbb{E}[\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})]\leq\mathbb{E}\Big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\Big],$
we obtain
$\mathbb{E}\Bigg[\int_{0}^{t}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg]$
$=\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})]\Big)\,\mathrm{d}s$
$\leq\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s.$
Therefore,
$C_{1}\,\mathbb{E}\Bigg[\Bigg(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg)^{1/2}\Bigg]$
$\quad\leq C_{1}\Big(\mathbb{E}[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]\Big)^{1/2}\Bigg(\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s\Bigg)^{1/2}.$
Applying Young’s inequality $uv\leq\tfrac{1}{2}\epsilon u^{2}+\tfrac{1}{2\epsilon}v^{2}$, let $\epsilon=\frac{1}{4C_{1}}$ we have
$C_{1}\,\mathbb{E}\Bigg[\Bigg(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg)^{1/2}\Bigg]$
$\quad\leq\frac{C_{1}\epsilon}{2}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\frac{C_{1}}{2\epsilon}\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s$
$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+2{C_{1}}^{2}\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s.$
Finally, using
$\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
$\quad\leq\max\{\tilde{L}_{1},\tilde{L}_{2}\}\Big(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big),$
we obtain
$C_{1}\,\mathbb{E}\Bigg[\Bigg(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg)^{1/2}\Bigg]$ (2.3)
$\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+2{C_{1}}^{2}\,\max\{\tilde{L}_{1},\tilde{L}_{2}\}$
$\quad\times\int_{0}^{t}\Big(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s.$
By using BDG’s inequality and (2.3), we have
$\mathbb{E}\left[\sup_{0\leq r\leq t}\left|\int_{0}^{r}\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\,\mathrm{d}W_{s}\rangle\right|\right]$ (2.4)
@@
$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+K_{1}\int_{0}^{t}\left(\mathbb{E}\Big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\mathbb{E}\Big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\Big]\right)\,\mathrm{d}s,$
where $K_{1}\;=\;2{C_{1}}^{2}\cdot\,\max\{\tilde{L}_{1},\tilde{L}_{2}\}$. In the same way for the term driven by $W$, we have
$\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$ (2.5)
$\quad\leq C_{2}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot|\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s})|^{2}\,\mathrm{d}s\right)^{1/2}\right]$
$\quad\leq C_{2}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot\left(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\right)\,\mathrm{d}s\right)^{1/2}\right]$
$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+K_{2}\int_{0}^{t}\Big(\mathbb{E}\big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\big]+\mathbb{E}\big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\big]\Big)\,\mathrm{d}s,$
where $K_{2}=2C_{2}^{2}\cdot\max\{\tilde{L}_{1},\tilde{L}_{2}\}$.
As for the jump term, let
$M_{r}:=\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z),\qquad r\in[0,t].$
Note that
$\Big|\big|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z\big|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big|$
$\quad\leq 2\,|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|\,\cdot\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|\,|z|+\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|^{2}\,|z|^{2}.$
By the Kunita’s inequality (see [14, Theorem 4.4.23]) for compensated Poisson integrals there exists a constant $C_{3}>0$ such that
$\mathbb{E}\Big[\sup_{0\leq r\leq t}|M_{r}|\Big]\leq C_{3}\Bigg\{\mathbb{E}\Big[\Big(\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\big\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},\;c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big\rangle\,z$
$+\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|^{2}\,|z|^{2}\Big|^{2}\,\nu(\mathrm{d}z)\,\mathrm{d}s\Big)^{1/2}\Big]+\;\mathbb{E}\Big[\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\big\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},\;c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big\rangle\,z$
$+\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|^{2}\,|z|^{2}\Big|\,\nu(\mathrm{d}z)\,\mathrm{d}s\Big]\Bigg\}.$
Using $|a+b|^{2}\leq 2|a|^{2}+2|b|^{2}$, $|a+b|\leq|a|+|b|$ and the Assumption (A7), we get
$\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\rangle z+|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s}|^{2}|z|^{2}\Big|^{2}\nu({\rm{d}}z){\rm{d}}s$
$\quad\leq\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}8|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\cdot|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})|^{2}|z|^{2}+2|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})|^{4}|z|^{4}\,\nu({\rm{d}}z){\rm{d}}s$
$\quad\leq\int_{0}^{t}8|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\big(\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big){\rm{d}}s$
$\qquad+2\big(\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big)^{2}\bigg(\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu(\mathrm{d}z)\bigg)^{-2}\int_{\mathbb{R}^{d}_{0}}|z|^{4}\nu({\rm{d}}z){\rm{d}}s,$
and
$\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\rangle z+|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s}|^{2}|z|^{2}\Big|\nu({\rm{d}}z){\rm{d}}s$
$\quad\leq\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}2|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|\cdot|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})||z|+|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s}|^{2}|z|^{2}\,\nu({\rm{d}}z){\rm{d}}s$
$\quad\leq\int_{0}^{t}2|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|\sqrt{\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})}\bigg(\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu(\mathrm{d}z)\bigg)^{-1/2}\int_{\mathbb{R}^{d}_{0}}|z|\nu({\rm{d}}z){\rm{d}}s$
$\qquad+\int_{0}^{t}(\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})){\rm{d}}s.$
Hence, similar to the term driven by $W$,
$\mathbb{E}\Big[\sup_{0\leq r\leq t}\Big|\int_{0}^{r}\int_{\mathbb{R}^{d}_{0}}\langle Y_{s}^{\mu}-Y_{s}^{\nu},\;c(Y_{s}^{\mu},\mu_{s},z)-c(Y_{s}^{\nu},\nu_{s},z)\rangle\,\tilde{N}(\mathrm{d}s,\mathrm{d}z)\Big|\Big]$ (2.6)
$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+K_{3}\int_{0}^{t}\Big(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s,$
where $K_{3}$ depends on $\tilde{L}_{1},\tilde{L}_{2}$ and the moments $\int_{\mathbb{R}^{d}_{0}}|z|\nu({\rm{d}}z),\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu({\rm{d}}z),\int_{\mathbb{R}^{d}_{0}}|z|^{4}\nu({\rm{d}}z)$. Substituting (2.4), (2.5), (2.6) into (2.2), we obtain